OpenStudy (anonymous):

What is the simplest form of the radical expression? 3^3√2a-6^3√2a

OpenStudy (anonymous):

\(3x-6x=-3x\)

OpenStudy (anonymous):

in your case \(x=\sqrt[3]{2a}\)

OpenStudy (anonymous):

so that's the answer? :)

OpenStudy (anonymous):

the answer is \(-3\sqrt[3]{2a}\) if i read the question correctly

OpenStudy (anonymous):

okay thank you :)

OpenStudy (anonymous):

yw